OpenStudy (anonymous):

The perimeter of a rectangle plot of land is 800 feet. The length is 20 feet less than twice the width. What is the length and width of the lot?

OpenStudy (anonymous):

assume that the width is x. 2(x)+2(2x-20)= 800 6x=840 x=140 lastly, width=x=140 feet length= 2x-20=260 feet correct me if im wrong

OpenStudy (anonymous):

that looks similar to what my teacher did in class im just not getting it

OpenStudy (anonymous):

how did you get 840?

OpenStudy (anonymous):

2(x)+2(2x-20)= 800 6x-40=800 6x=840 x=140

OpenStudy (anonymous):

your combining like terms?

OpenStudy (anonymous):

yes. it would be easier if you draw a rectangle and mark each side with each function stated above.